### v3

# DiagramQG: Concept-Focused Diagram Question Generation via Hierarchical Knowledge Integration

Abstract
Visual Question Generation (VQG) has gained significant attention due to its potential in educational applications. However, VQG research mainly focuses on natural images, largely neglecting diagrams in educational materials used to assess students’ conceptual understanding. To address this gap, we construct DiagramQG, a dataset containing 8,372 diagrams and 19,475 questions across various subjects. DiagramQG introduces concept and target text constraints, guiding the model to generate concept-focused questions for educational purposes. Meanwhile, we present the Hierarchical Knowledge Integration framework for Diagram Question Generation (HKI-DQG) as a strong baseline. This framework obtains multi-scale patches of diagrams and acquires knowledge using a visual language model with frozen parameters. It then integrates knowledge, text constraints, and patches to generate concept-focused questions. We evaluate the performance of existing VQG models, open-source and closed-source vision-language models, and HKI-DQG on the DiagramQG dataset. Our novel HKI-DQG consistently outperforms existing methods, demonstrating that it serves as a strong baseline. Furthermore, we apply HKI-DQG to four other VQG datasets of natural images, namely VQG-COCO, K-VQG, OK-VQA, and A-OKVQA, achieving state-of-the-art performance.
1 Introduction
Visual Question Generation (VQG) is a crucial research area in computer vision and natural language processing. It aims to automatically generate meaningful questions from input images. The evolution of VQG approaches has witnessed a significant shift from traditional rule-based and template-based methodologies to sophisticated neural network architectures, encompassing both unconstrained [26] and constrained [13, 11, 34, 33] paradigms. While early VQG research focused primarily on generating questions using only image inputs [26], contemporary research has expanded to incorporate various guiding information to enhance question generation specificity, including answers [13, 15], answer types [11, 35], image region [34], question types [7], and knowledge triple [33], thereby significantly improving the quality and relevance of generated questions. We demonstrate various existing text constraints as inputs for visual question generation, as shown in (a) of Figure 1.
Figure 1: Comparison of different types of text constraints in visual question generation on natural images (a) and on the DiagramQG dataset (b).
However, current VQG research focuses on natural images, neglecting diagrams commonly popular in educational materials. This fails to meet educational requirements for assessment. Diagrams play a vital role in education by presenting information in a structured engaging manner, enabling students to understand complex concepts more effectively [4]. Moreover, these concept-focused questions based on diagrams assess students’ ability to explain, analyze, and apply the information presented in diagrams using concepts [12] on diverse subjects. Therefore, generating questions based on diagrams is crucial for assessing students’ ability to apply the learned concepts in a practical context.
To bridge the gap between current research and educational requirements, we have constructed DiagramQG, the first dataset containing 8,372 diagrams and 19,475 questions, covering 4 subjects, 15 courses, and 169 concepts, by integrating existing diagram question-answering datasets [10, 39, 48, 22] with curriculum materials we collected. These questions provide a more comprehensive evaluation of student’s understanding of the concepts across different subjects. Meanwhile, the DiagramQG dataset introduces concept and target text constraints, guiding the model to generate questions that assess students’ understanding of concepts and include the target text, as shown in (b) of Figure 1. The target text serves as guidance to control question diversity, as each diagram can generate multiple concept-focused questions. Related research shows that answer-based constraints tend to generate context-independent questions [17]. Question-type constraints often deviate from the intended assessment objectives [33]. Moreover, knowledge triple-based constraints frequently result in mere triple expansions rather than meaningful questions [24]. This highlights the need for novel text constraints to generate educationally meaningful questions.
Compared to natural images, diagrams require a significant amount of subject-specific background knowledge for understanding [40], and they typically contain a higher number of objects [9]. Thus, tackling the Diagram Question Generation (DQG) task requires a framework that can incorporate subject-specific knowledge, comprehend object relationships within diagrams, and generate meaningful questions accordingly. We propose a novel Hierarchical Knowledge Integration framework for Diagram Question Generation (HKI-DQG). The framework leverages CLIP to identify relevant multi-scale diagram patches, from which vision-language models like BLIP and Qwen2-VL extract knowledge. After T5 filters the extracted knowledge to ensure relevance to text constraints, the framework integrates these filtered insights with text constraints and multi-scale diagram patches to generate questions. Notably, this approach doesn’t require training the vision-language model, which improves scalability and reduces costs. We evaluate HKI-DQG against current VQG and vision-language models on our DiagramQG dataset. Additionally, we conduct experiments with HKI-DQG on four natural images of VQG datasets, validating HKI-DQG’s generalizability and effectiveness. Our contributions are summarized as follows:
  • We introduce the comprehensive DiagramQG dataset, the first question generation dataset specifically designed for diagrams, and develop a novel text constraint that guides the model in generating concept-focused questions.
  • We propose an adaptable framework based on different frozen-parameter vision-language models, addressing DQG tasks at a lower cost, and conduct experiments on both DiagramQG and four natural image VQG datasets.
  • We thoroughly evaluate the performance of current mainstream VQG models, various open-source and closed-source vision-language models, and our proposed framework on the newly constructed DiagramQG dataset.
2 Related Work
2.1 Visual Question Generation
VQG task has progressed from rule-based and template-based methods to advanced neural network architectures, including both unconstrained and constrained approaches. The unconstrained approach often produces generic questions with insufficient image specificity, restricting the development of unconstrained VQG methods [1, 26]. Researchers have explored various text constraints to improve question generation quality, including answer-guided methods [11, 42, 45, 19], question types [7], answer categories [37], and knowledge-enhanced approaches [43, 2]. Recent works have proposed cross-topic VQG models [18] and dual contrastive learning approaches [24] to address unseen topics and improve visual-textual alignment. However, current methods face challenges in balancing question diversity and effective utilization of visual information, emphasizing the urgent need for innovative text constraints and robust methods that ensure constraint validity, question naturalness, and diversity in the use of visual information.
2.2 Diagram Question Answering
Diagram Question Answering (DQA) requires higher reasoning ability and background knowledge compared to Visual Question Answering (VQA). Previous research has focused on enhancing diagram understanding through pre-training approaches [8, 23, 44]. Additionally, studies investigate the use of prompting methods with Large Language Models to harness their extensive prior knowledge for DQA [22, 49, 46, 40]. Aligning the information in the diagram with background knowledge is crucial for understanding the diagram and achieving effective DQA performance.
Figure 2: Four different examples of different subjects in DiagramQG dataset.
Figure 3: Domain diversity in DiagramQG. Each color corresponds to one subject: Natural Science (blue), Earth Science (yellow), Applied Science (green), and Social Science (orange).
3 The DiagramQG Dataset
The DiagramQG is a dataset of questions covering 4 subjects, 15 courses, and 169 concepts, consisting of 19,475 unique questions and 8,372 diagrams. These components form 44,472 combinations of tuples (target & concept text constraint, diagram, question). We provide examples from four subjects in the DiagramQG dataset, as shown in Figure 2. The diagram list annotates objects appearing in the diagram, which will be used in later evaluation metrics.
3.1 Data Collection
The data collection process can be divided into four distinct phases. This process begins by gathering content from well-established datasets [10, 39, 48, 22], supplemented by searching for relevant images on Hugging Face, GitHub, and Roboflow, with all data following CC BY-SA 4.0, CC BY-NC-SA, and MIT licenses. This initial phase yields a substantial raw dataset containing over 25,000 diagrams and 60,000 questions spanning multiple subjects.
Subsequently, the collected diagram and questions are organized into four primary subjects and further categorized into 15 specific courses. This structuring process involves mapping the questions to their corresponding concepts, resulting in the identification of 169 distinct concepts.
Next, experienced teachers from different academic subjects with teaching experience annotate target text constraints for each diagram-question pair under their specialized concepts. This detailed annotation effort produces more than 70,000 unique combinations, as some questions are associated with different target & concept text constraints, such as the same concept but different targets.
Finally, another group of subject teachers screens all current combinations based on their educational utility, using a 100-point scale with a passing threshold of 60 points. The resulting DiagramQG dataset contains 8,372 diagrams, 19,475 questions, and 44,472 complete combinations, sourced from ScienceQA (34.1%), TQA (24.3%), CSDQA (10.8%), ADE (9.2%), and our collection (21.6%).
3.2 Data Analyse
3.2.1 Domain & Question diversity.
Num. of Q. Num. of I. Object on I. Images Text Constraint Science subjects
VQAv2.0 1.1M 20k 3.5 natural answer N/A
FVQA 5,826 2k 2.9 natural answer N/A
VQG-COCO 25,000 5k 3.3 natural image, caption N/A
K-VQG 16,098 13K 2.7 natural knowledge triple N/A
ScienceQA 21,208 10,332 4.3 diagram answer natural, social
CSDQA 3,494 1,294 5.2 diagram answer computer science
TQA 26,260 3,455 6.8 diagram answer natural, earth
ADE 3.945 3.945 7.2 diagram answer natural
DiagramQG 19,475 8,372 11.2 diagram target, concept natural, earth, application, social
Table 1: Comparing core characteristics of major VQG datasets, DQA datasets and DiagramQG, where Q. and I. mean question and image.
Figure 4: Question distribution in DiagramQG.
Figure 5: Distribution of diagrams, questions, and questions per diagram ratios across different concepts in DiagramQG.
The DiagramQG dataset encompasses a broad spectrum of educational domains, categorized into four primary subjects: Natural Science (blue), Earth Science (yellow), Applied Science (green), and Social Science (orange), as shown in Figure 3. The dataset is structured hierarchically, with samples first classified by subject, then further divided into specific courses (e.g., Biology), and ultimately categorized into concepts (e.g., Ecological interactions). Figure 3 presents a treemap visualization illustrating the dataset’s breadth and depth, encompassing 4 subjects, 15 courses, and 169 concepts across several academic domains.
The DiagramQG dataset is distinguished by its diverse range of scientific questions, which exhibit significant variation in both question formulation and complexity. As clearly depicted in Figure 4, an extensive analysis of the first four words in question texts reveals a wide array of question types and structures. The remarkable diversity is further exemplified by the variation in question length, ranging from 4 to 50 words, with an average length of 9.52 words, underscoring the comprehensive nature of the dataset.
An in-depth analysis of the distribution of diagrams and questions across different concepts reveals a distinct long-tail distribution pattern. As shown in Figure 5, the ratio of questions per diagram for each concept also exhibits a pronounced long-tail phenomenon. This skewed distribution indicates that while certain concepts are well-represented, a considerable number of concepts have limited coverage—a common challenge in educational datasets, representative of a typical occurrence in educational environments.
3.2.2 Comparisons to Other Datasets.
Table 1 presents a comprehensive comparison between the DiagramQG dataset and existing VQG and DQA datasets. The comparative analysis reveals that DiagramQG demonstrates greater educational utility compared to VQG datasets, which primarily focus on general image description and common-sense reasoning tasks, and surpasses existing DQA datasets in subject matter coverage. Moreover, DiagramQG stands out as the first large-scale multimodal scientific question generation dataset that incorporates non-answer words or phrases and concepts across different subjects as text constraints. A notable aspect of our analysis involves the examination of object density per diagram across the nine datasets. This is particularly significant considering that diagrams inherently contain more complex and information-rich content compared to natural images.
3.2.3 Challenges in DiagramQG dataset
Based on our comprehensive comparative analysis, DiagramQG presents three unique challenges that differentiate it from existing visual question generation datasets:
  • Domain-specific knowledge requirement: Unlike other existing VQG datasets that rely on general common sense, DiagramQG consistently encourages models to possess and apply different concepts across various subjects to generate meaningful and practical questions.
  • Long-tail distribution phenomenon: The inherent complex long-tail distribution in DiagramQG, where concept coverage ranges from abundant to limited, challenges models to generalize and optimize performance across both well-represented and underrepresented concepts.
  • High object information density: The significantly high density of object information in the diagrams complicates content interpretation and risks overlooking critical details, demanding models to possess capabilities in capturing and processing complex visual information.
4 Baseline
4.1 Problem Definition
The Diagram Question Generation (DQG) task aims to generate a concept-focused question $q$ given a diagram $d$, a target text $t$, and a concept text $c$. The generated question should contain the target text $t$ and effectively assess students’ understanding of the specified concept $c$. This task can be formulated as a conditional generation problem, represented as $p(q|d,t,c)$, where a multimodal model maps visual and textual information into a joint embedding space before decoding questions that satisfy both the text constraint and the conceptual assessment requirement.
Figure 6: (a) Process the diagram into a pyramid structure of patches. (b) Our DiagramQA baseline, HKI-DQG, consists of three distinct stages: HierKnowExtract, KnowSelect, and KnowFusionQG. In this framework, orange modules indicate trainable parameters, blue modules represent fixed parameters, and gray modules denote components without learnable parameters.
4.2 Architecture
We propose the Hierarchical Knowledge Integration (HKI-DQG) framework as a baseline for the DQG task. This framework is designed to be compatible with any vision-language foundation model and implements a three-stage pipeline for question generation: HierKnowExtract, KnowSelect, and KnowFusionQG, as shown in Figure 6. The HierKnowExtract stage extracts knowledge $K_{s}$ from multi-scale diagram patches $P_{s}$ using vision-language models with frozen parameters. The KnowSelect stage selects the $m$ most relevant knowledge sentences $K_{(t,c)}$ based on text $t$ and concept $c$. The KnowFusionQG stage integrates $t$, $c$, $P_{d}$, and $K_{(t,c)}$ to generate the final question $\hat{q}$.
4.2.1 HierKnowExtract
The HierKnowExtract stage obtains diagram patches $P_{s}$ of different scales related to the target & concept text constraints and uses vision-language models to extract the knowledge $K_{s}$ contained in all patches. This process begins with a hierarchical decomposition of diagram $d$ into ordered patches $P_{d}$ across an $n$-layered pyramid structure, followed by visual encoding using the CLIP Image Encoder [31] to get $F^{l}$, as formulated in Equation 1.
{$&F_{d}=\{F^{l}\}_{l=1}^{n}\\ &P_{d}=\{p^{l}_{i,j}\mid l\in[1,n],i,j\in[1,l]\}\\ &F^{l}=\{f^{l}_{i,j}=\text{CLIP}(p^{l}_{i,j})\mid p^{l}_{i,j}\in P_{d}\}\\ &p^{l}_{i,j}=d\left[\frac{i-1}{l}H:\frac{i}{l}H,\frac{j-1}{l}W:\frac{j}{l}W\right]\\ \end{aligned}\right.$ (1)
where $H$ and $W$ denote the height and width of the input diagram, respectively.
For textual components, the T5 Encoder [32] is employed to process both the target text $t$ and concept text $c$ independently. We introduce a learnable linear transformation $W_{h}$ to ensure compatibility between the CLIP Image Encoder [31] and T5 Encoder [32] feature spaces. This effectively facilitates unified similarity computation and subsequent operations in the KnowFusionQG phase. The process ultimately culminates in the selection of the most semantically relevant patch from each hierarchical layer to form the set $P_{s}$, as shown in Equation 2.
{$&e_{t}=\text{T5}_{\text{enc}}(t),\quad e_{c}=\text{T5}_{\text{enc}}(c)\\ &s^{l}_{i,j}=\text{sim}(W_{h}f^{l}_{i,j},e_{t})+\text{sim}(W_{h}f^{l}_{i,j},e_{c})\\ &P_{s}=\{p^{l}_{i^{*},j^{*}}\mid(i^{*},j^{*})=\operatorname*{argmax}_{i,j}s^{l}_{i,j},l\in[1,n]\}\end{aligned}\right.$ (2)
Selected patches $P_{s}$, target text $t$, and concept text $c$ are carefully fed into a large-scale vision-language model (VLM) like Qwen2-VL or BLIP to obtain diverse patch-specific knowledge. The resulting knowledge set $K_{s}$ is systematically generated according to Equation 3.
{$&K_{s}=\{k^{l}\mid l\in[1,n]\}\\ &k^{l}=\text{VLM}(p^{l}_{i^{*},j^{*}},t,c),\quad\forall p^{l}_{i^{*},j^{*}}\in P_{s}\end{aligned}\right.$ (3)
where $k^{l}$ is the text paragraph retrieved by the VLM, containing several knowledge sentences. To optimize the learning process, the CLIP Image Encoder [31], T5 Encoder [32], and the linear transformation $W_{h}$ employed in this phase share parameters with their counterparts in the third phase, where gradient updates are effectively propagated.
4.2.2 KnowSelect
The KnowSelect stage selects the $m$ most relevant knowledge sentences $K_{(t,c)}$ from the extensive knowledge set $K_{s}$ based on target text $t$ and concept text $c$. The process begins with encoding the knowledge set $K_{s}$ and the text constraints using the T5 Encoder [32], as formulated in Equation 4.
{$H_{K}&=\text{T5}_{\text{enc}}(K_{s})\\ H_{t,c}&=\text{T5}_{\text{enc}}(\text{Prompt}(t,c))\end{aligned}\right.$ (4)
where $\text{Prompt}(t,c)$ synthesizes $t$ and $c$ into a coherent prompt following the template: Given the target text $t$, identify key knowledge related to the concept $c$.
To quantify the semantic relevance between the knowledge set ($K_{(t,c)}$) and text constraints ($t$ and $c$), we employ a scaled dot-product attention mechanism. This process computes attention scores between the knowledge tokens and the text constraint, followed by knowledge selection, as formulated in Equation 5:
{$&A=\text{softmax}\left(\frac{H_{K}H_{t,c}^{T}}{\sqrt{d_{k}}}\right)\\ &K_{(t,c)}=\text{top-m}(AH_{K})\end{aligned}\right.$ (5)
where $d_{k}$ denotes the dimensionality of the hidden representations. The top-m() operator selects the $m$ most semantically relevant knowledge sentences based on attention scores, creating a knowledge set $K_{(t,c)}$ that best matches the text constraints.
4.2.3 KnowFusionQG
The KnowFusionQG stage integrates the selected diagram patches $P_{s}$, target text $t$, concept text $c$, and knowledge set $K_{(t,c)}$ to generate the concept-focused question through a multimodal fusion mechanism. This process begins with encoding the visual and textual inputs into language and vision representations, as shown in Equation 6.
{$&H_{t}=\text{T5}_{\text{enc}}(\text{Prompt}[t;c;K_{(t,c)}])\\ &H_{v}=W_{h}\cdot\text{CLIP}(P_{s})\end{aligned}\right.$ (6)
where $W_{h}$ is also used in HierKnowExtract phase. This $\text{Prompt}[t;c;K_{(t,c)}]$ synthesizes $t$, $c$ and $K_{(t,c)}$ into a coherent prompt following the template: Generate the question including Target: $t$ to assess Concept: $c$ with the knowledge: $K_{(t,c)}$. To capture the intricate relationships between textual and visual representations, a cross-modal attention mechanism followed by a gated fusion network is employed, as formulated in Equation 7.
{$&H^{attn}_{v}=\text{Softmax}(\frac{H_{t}H_{v}^{T}}{\sqrt{d_{k}}})H_{v}\\ &\lambda=W_{t}H_{t}+W_{v}H^{attn}_{v}\\ &H_{fuse}=H_{t}+\tanh(\lambda)\cdot H^{attn}_{v}\end{aligned}\right.$ (7)
Finally, the fused output $H_{fuse}$ is fed into the T5 decoder [32] to predict the target question $\hat{q}$. This integration of visual and text information through cross-modal attention and gated fusion enables the model to generate questions that are contextually relevant and conceptually focused.
5 Experiments
5.1 Implementation Details
The experimental framework uses T5 Base and Large architectures along with the CLIP (ViT-B/32) model. Optimization is performed with the AdamW optimizer, applying a learning rate of 1e-5 for CLIP and 5e-5 for other components, running through 20 epochs of fine-tuning. Key parameters include a maximum input sequence length of 256, an output sequence length of 32, and a batch size of 32, with experiments conducted on one NVIDIA A800 80G GPU. The DiagramQG dataset is split into training, validation, and testing sets with a ratio of 70:5:25, ensuring no overlap of diagrams and questions across these sets. Additionally, each concept is represented in both validation and testing sets, despite some concepts having relatively few samples.
Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 Bert-Score METEOR CIDEr ROUGE-L In-Context Learning InternVL2 (26B) [3] 27.84 16.54 10.76 7.64 78.72 15.56 0.37 24.12 InternVL2 (40B) [3] 28.40 17.78 11.85 8.63 79.72 16.91 0.39 25.35 MiniCPM-V2(8B) [47] 24.25 12.51 7.86 4.89 74.64 11.49 0.28 21.96 DeepSeek-VL(7B) [20] 23.96 14.14 8.53 5.51 76.12 14.96 0.30 23.42 Qwen2-VL(2B) [38] 26.87 15.53 9.35 5.59 75.16 12.76 0.29 23.21 Qwen2-VL(7B) [38] 28.64 17.71 11.51 8.59 76.68 15.18 0.45 24.79 Qwen2-VL(72B) [38] 35.82 25.17 18.94 15.21 82.19 19.84 0.71 31.76 GLM4-V pro 31.89 20.05 13.87 10.68 81.16 16.52 0.62 27.41 Claude-3.5-Sonnet (T) 22.76 13.01 8.05 5.16 78.77 13.26 0.45 20.51 GPT-4o (T) 19.31 10.22 6.17 3.65 80.54 11.02 0.36 20.7 Claude-3.5-Sonnet 41.86 35.15 30.14 26.20 83.54 30.39 1.78 39.67 GPT-4o 43.72 37.22 32.26 28.13 84.16 31.50 1.91 43.39 Fine-tuned K-VQG [33] 22.37 16.97 13.51 11.05 73.88 27.48 0.45 28.55 Patch-TRM [21] 23.02 17.63 14.08 11.67 74.68 28.74 0.47 30.50 ConVQG (BLIP) [24] 31.21 23.42 18.47 15.06 74.71 29.89 0.56 31.24 KC-VQG (BLIP) [17] 33.13 26.56 21.75 17.52 79.84 31.59 1.03 34.26 KC-VQG (Q-2B) [17] 36.11 28.74 23.59 19.72 83.54 33.93 1.32 37.54 Ours HKI-DQG(BLIP, T5-B) 47.98 40.25 35.31 31.87 84.27 41.68 2.91 46.31 HKI-DQG(Q-2B, T5-B) 58.76 52.64 48.52 46.94 88.74 52.44 4.15 58.21 HKI-DQG(Q-7B, T5-B) 60.62 54.68 50.54 47.48 89.52 53.09 4.19 59.97 HKI-DQG(BLIP, T5-L) 50.63 43.61 38.94 35.69 85.44 43.84 3.11 49.75 HKI-DQG(Q-2B, T5-L) 60.03 54.14 50.07 47.05 89.28 53.21 4.21 59.62 HKI-DQG(Q-7B, T5-L) 61.63 55.62 51.46 48.41 89.81 54.26 4.32 60.69
Table 2: Results on DiagramQG, where GPT-4o (T) and Claude-3.5-Sonnet (T) are meant to generate questions using only text constraints, where, Q-2B, Q-7B, T5-B, T5-L mean, Qwen2-VL-2B, Qwen2-VL-7B, T5-Base and T5-Large.
[…]
Table 3: Evaluation results on K-VQG dataset.
5.2 Baselines
VQG approaches can be classified into two main categories: Fine-tuned and In-Context Learning methods. Fine-tuned methods use models like BERT [6], T5 [32], or GPT-2 [30] for question generation, with examples including IM-VQG [11], K-VQG [33], Patch-TRM [21], ConVQG [24], LV2-Net [19], and KC-VQG [17]. In-Context Learning methods utilize large vision-language models to extract relevant information from diagrams and generate questions using three reference questions. This includes open-source models like Qwen2-VL [38], MiniCPM-V [47], DeepSeek-VL [20], and InternVL2 [3], as well as closed-source models such as GLM4-V pro, Claude-3.5 Sonnect, and GPT-4o. Models such as Llava 1.5 [16], and CogVLM [41] do not perform satisfactorily on the DiagramQG dataset, likely caused by limited exposure to diagram datasets.
Figure 7: The solid line represents concept-wise BLEU scores, while the dotted line indicates the mean performance.
5.3 Evaluation metrics
We evaluate our model using established language generation metrics, including BLEU [27], Bert Score, METEOR [5], CIDEr [36], and ROUGE [14]. We evaluate model performance using the pycocoevalcap package to measure alignment between generated and ground truth questions.
5.4 Results
We evaluate various models on our DiagramQG dataset, as shown in Table 2. Among the in-context learning methods that utilize Vision-Language Models (VLMs), models with larger parameter counts generally perform better, indicating their abundant knowledge across different subjects. Interestingly, Qwen2-VL 2B achieves impressive and competitive results despite having fewer parameters, making it a suitable base model for further comprehensive experiments. Tests conducted with GPT-4o and Claude-3.5-Sonnet using only text inputs result in reduced generation performance, highlighting the critical importance of incorporating diagrams in the question generation process. Among the fine-tuned methods, KVQG and Patch-TRM do not leverage VLMs for additional knowledge, while ConVQG and KC-VQG do. The latter models demonstrate superior performance, with KC-VQG showing improved results as VLM parameters scale up. HKI-DQG framework achieves the best performance across all quantitative evaluation metrics, regardless of whether it is based on T5-base or T5-large, particularly when using the larger-scale Qwen2-VL model. These impressive results validate HKI-DQG as an effective baseline for future research by demonstrating the value of incorporating subject knowledge in the DQG task. We further analyze the performance of HKI-DQG across different concepts, as illustrated in Figure 7. The results exhibit a long-tail distribution, where concepts with more training samples consistently and significantly perform better.
[…]
Table 4: Evaluation results on VQG-COCO, OK-VQA and A-OKVQA datasets.
5.5 Results on Other Dataset
We conduct comprehensive evaluations of HKI-DQG’s generalization capabilities across four prominent VQG datasets of natural images: VQG-COCO, OK-VQA, A-OKVQA, and K-VQG, with results presented in Tables 4 and 3. The experimental results demonstrate that HKI-DQG consistently outperforms existing approaches across most evaluation metrics, with notably superior performance in CIDEr and METEOR scores, suggesting enhanced semantic comprehension and contextually coherent question generation.
H.K.E K.S K.F BLEU-4 METEOR CIDEr
$\checkmark$ $\checkmark$ $\checkmark$ 46.94 52.44 4.15
$\checkmark$ $\checkmark$ 44.65 50.07 3.86
$\checkmark$ $\checkmark$ 40.87 48.53 3.27
$\checkmark$ 39.13 46.87 3.13
$\checkmark$ 16.47 28.24 0.53
Table 5: Results on DiagramQG with Qwen2-VL 2B & T5-Base to obtain the results, where H.K.E, K.S and K.F mean HierKnowExtract, KnowSelect and KnowFusionQG.
5.6 Ablation experiment
Given that KnowSelect is inherently based on HierKnowExtract, isolated evaluation of KnowSelect is not feasible. The ablation study results are presented in Table 5. The findings demonstrate that incorporating relevant knowledge substantially improves question generation quality. Notably, KnowSelect shows greater impact than KnowFusionQG, indicating that while HierKnowExtract generates abundant knowledge, KnowSelect effectively filters and retains the most relevant information.
Setting BLEU-4 METEOR CIDEr
n=1,m=4 40.69 45.93 3.31
n=2,m=4 43.76 50.68 3.64
n=4,m=4 46.12 51.95 4.02
n=3,m=1 42.63 48.81 3.72
n=3,m=2 44.59 50.26 3.88
n=3,m=3 45.65 51.70 4.05
n=3,m=5 46.69 52.28 4.12
n=3,m=4 46.94 52.44 4.15
Table 6: Results on DiagramQG under different settings. We employ HKI-DQG (Qwen2-VL 2B & T5-Base) to obtain the results.
5.7 Parameter Sensitive Study
Our experiments involve two manually set key parameters: dividing the diagram into an $n$-layer pyramid of patches and selecting $m$ sentences of background knowledge for question generation. Based on our analysis, Table 6 clearly shows the detailed results with $n$ ranging from $1$ to $4$, with the model performance consistently peaking at $n=3$, strongly suggesting that overly fine diagram division leads to noisy knowledge extraction by VLM, significantly affecting model performance. Regarding the second important parameter, Table 6 systematically shows experimental results for m ranging from $1$ to $5$, definitively peaking at $m=4$. While excessive background knowledge notably affects the performance of HKI-DQG, the relatively minor overall impact clearly demonstrates the framework’s impressive resilience to redundant input information.
Figure 8: Comparison between HKI-DQG and GPT-4o using two distinct examples from the DiagramQG dataset.
5.8 Case Study
Figure 8 illustrates examples from the DiagramQG dataset comparing the HKI-DQG framework and GPT-4o, showcasing HKI-DQG’s superior question generation capabilities. While advanced models like GPT-4o can generate coherent and fluent question texts, those questions often fail to effectively test students’ deeper understanding of concepts. Compared to GPT-4o, HKI-DQG consistently demonstrates significant advantages in generating questions with stronger concept relevance. For a more detailed case study analysis and results, please refer to the supplementary materials.
6 Conclusion
This study introduces DiagramQG, the first dataset specifically designed for diagram question generation. We propose HKI-DQG, a novel framework that leverages hierarchical knowledge integration to generate concept-focused questions. Experiments show that HKI-DQG framework outperforms existing VQG and popular vision-language models. This work advances intelligent education [25] by enabling automated question generation for personalized learning.
Supplementary Material
In this supplementary material, we provide more details on our implementation and experiments as follows:
  • Section A: More details on implementation;
  • Section B: More details on baselines;
  • Section C: More details on DiagramQG;
  • Section D: More experiments on other VQG datasets;
  • Section E: More case studies on Diagram.
A. More details on implementation
In our experimental setup, we developed a framework based on the T5 architecture, utilizing both its Base and Large variants with 220M and 770M parameters respectively, alongside CLIP (ViT-B/32) for visual understanding capabilities. The framework’s trainable parameters, excluding pre-trained components, were initialized following a normal distribution ($\mu=0$, $\sigma=0.02$). For optimization, we implemented the AdamW optimizer with a dual learning rate strategy: a conservative $51e-$ 5 for CLIP components to maintain visual feature integrity, and a higher $55e-$ 5 for remaining components. The training process spanned $20$ epochs, incorporating linear learning rate warmup during the initial $2$ epochs followed by cosine decay. Our implementation featured specific configurations including $256$ tokens for maximum input sequence length, $32$ tokens for output sequence length, a batch size of $32$, $4$ gradient accumulation steps, $0.01$ weight decay, and $0.1$ dropout rate, with FP16 mixed precision training enabled. All experimental procedures were executed on a hardware setup consisting of two NVIDIA Tesla A100 80G GPUs, utilizing CUDA 11.8 and PyTorch 1.13.1.
B. More details on baselines
Visual Question Generation (VQG) approaches can be classified into two main categories: fine-tuning-based methods and large vision-language models.
For Fine-tuning-based methods, using the same data split approach as our method, the dataset is divided into training, validation, and test sets in a ratio of 7:1:2. Models are trained on the training set, validated on the validation set, and final results are reported on the test set.
K-VQG [33] integrates UNITER, a multi-modal transformer, to encode both visual features (extracted via Faster R-CNN) and masked knowledge triplets. The model processes visual region features with positional information and combines them with tokenized knowledge triplets through a BART-based decoder, generating questions
Patch-TRM [21] employs a hierarchical patch-based transformer that processes diagrams by decomposing them into meaningful patches through a pyramid layout. The model combines ResNet and vision Transformer for visual processing, using attention mechanisms to fuse visual and textual features, enabling effective capture of both local details and global relationships for diagram-focused question generation.
ConVQG [24] introduces a contrastive learning framework with dual modality-specific objectives for visual question generation. By contrasting image-specific and text-guided features, the model generates diverse questions that are both image-specific and controllable through flexible textual constraints such as answers, captions, or knowledge triplets, enabling precise control over question content while maintaining visual grounding.
KC-VQG [17] presents a knowledge-guided framework that combines topic-aware visual attention and Large Language Model (LLM) generated knowledge for question generation. The model integrates three components: a topic-aware visual feature extractor, a knowledge extractor with discriminative filtering, and a GPT-2 based decoder, enabling it to generate questions that incorporate both explicit visual information and implicit commonsense knowledge about specified topics.
For large vision-language models (VLMs), whether open-source or closed-source, we selected three reference questions for each concept. These reference questions, along with target and concept textual constraints, are incorporated into the prompt during the VLM’s question generation process to obtain final outputs.
Open-Source Large vision-language models:
Qwen2-VL [38] introduces two key architectural innovations in vision-language modeling: a Naive Dynamic Resolution mechanism that processes arbitrary image resolutions by mapping them to dynamic visual tokens, and a Multimodal Rotary Position Embedding (M-ROPE) that decomposes positional information to capture 1D textual, 2D visual, and 3D video dimensions. Available in 2B, 7B, and 72B parameter versions, it employs a streaming architecture that enables efficient processing of varying input formats while maintaining strong performance on diagram-related tasks.
MiniCPM-V [47] presents a compact yet effective vision-language model that combines SigLip-400M visual encoder with Qwen2-7B language model, totaling 8B parameters. The model offers efficient multi-image and video understanding capabilities while maintaining competitive performance on vision-language tasks through its streamlined architecture and parameter-efficient design.
Statistic Number
Total unique diagrams 8,372
Total unique questions 19,475
Total combinations 44,472
Total Subject 4
Total Course 15
Total Concept 169
Train unique diagrams 5,610
Train unique questions 13,271
Train combinations 30,531
Val unique diagrams 922
Val unique questions 2,015
Val combinations 4,350
Test unique diagrams 1,839
Test unique questions 4,189
Test combinations 9,591
Table 7: Main statistics in DiagramQG
DeepSeek-VL [20] introduces a balanced approach to vision-language modeling that builds upon the DeepSeek language model series. The model emphasizes maintaining strong language capabilities while developing visual understanding, featuring high-resolution processing capabilities and a carefully curated training strategy. Through its systematic scaling methodology from 1B to 7B parameters, it achieves competitive performance in practical applications while maintaining efficiency in multi-modal processing.
InternVL2 [3] introduces a progressive alignment training strategy that scales efficiently from small to large language models while supporting diverse input modalities (text, images, video, and medical data). Built upon InternVL 1.5, the model incorporates an expanded training dataset covering multiple tasks (captioning, VQA, detection, grounding, and OCR) and employs VisionLLMv2 technology to enable various output formats including images, bounding boxes, and masks. This comprehensive approach allows the model to handle hundreds of vision-language tasks while maintaining expert-level performance.
Closed-Source Large vision-language models:
GLM4-V is a large-scale multimodal language model developed by ZhiPu, featuring outstanding visual understanding and language generation capabilities. It employs a unified pre-training framework, capable of handling multiple modalities such as text, images, and audio. GLM4-V demonstrates strong performance in tasks like visual question answering, image description, and cross-modal reasoning, and can quickly adapt to new scenarios through few-shot learning. The model supports both Chinese and English, excelling in multilingual understanding and generation.
Claude 3.5 Sonnet is a next-generation multimodal assistant developed by Anthropic, exhibiting exceptional performance in both visual and language understanding. It utilizes an innovative neural network architecture that allows for deep comprehension of image content and complex reasoning. The model has strict controls in terms of safety and ethics, capable of identifying and filtering inappropriate content. A notable feature is its strong contextual understanding and coherent conversational abilities.
GPT-4o is the latest large language model developed by OpenAI, possessing powerful multimodal understanding and generation capabilities. It can process image and text inputs and perform complex reasoning and knowledge integration. This model showcases remarkable few-shot learning abilities, quickly mastering new tasks with only a few examples. GPT-4o also excels in creativity, generating high-quality text, code, and other creative content.
Other Large vision-language models:
Llava 1.5 [16], and CogVLM [41], despite their strong performance in general vision-language tasks, showed limitations in diagram-specific question generation, possibly due to insufficient training on diagram-focused datasets.
Figure 9: Three different cases (Text Constraint Lost) in DiagramQG dataset.
Figure 10: Three different cases (Diagram Interpretation Bias) in DiagramQG dataset.
Figure 11: Three different cases (Concept Ambiguity) in DiagramQG dataset.
C. More details on DiagramQG
The DiagramQG dataset is a comprehensive collection of questions covering 4 subjects, 15 courses, and 169 concepts, consisting of 19,077 unique questions and 8,372 diagrams. These components form 44,074 combinations of (target & concept text constraint, diagram, question). This dataset aims to encourage models to generate questions that assess students’ conceptual understanding by leveraging the provided target & concept text constraints and diagrams, as shown in Figure 2. Accomplishing this task requires two key capabilities from computational models. First, they must demonstrate sophisticated visual comprehension abilities. Second, they need comprehensive domain knowledge across multiple academic disciplines. Following established methodological practices in machine learning, we implemented a stratified partitioning of the dataset, allocating 70% for training, 10% for validation, and 20% for testing purposes. The comprehensive statistical distribution across these partitions, along with the aggregate dataset metrics, is presented in Table 7.
Figure 12: Three different cases in DiagramQG dataset.
Figure 13: Three different cases in DiagramQG dataset.
D. More experimental details on other VQG datasets
Method BLEU-1 METEOR CIDEr
MDN [29] 36.0 23.4 0.51
MC-BMN [28] 40.7 22.6 0.50
ConVQG [24] 50.2 26.4 0.56
[…]
Table 8: Results on VQG-COCO, using captions as text constraint. We report BLEU-1 instead of BLEU-4 to be consistent with the comparison methods.
Method BLEU-4 METEOR CIDEr
IM-VQG [11] 11.44 17.07 0.26
K-VQG [33] 18.84 22.79 1.31
ConVQG [24] 20.01 22.66 1.53
[…]
Table 9: Results on the K-VQG dataset. The results of IM-VQG are reproduced based on the official code.
Method BLEU-4 METEOR CIDEr
IM-VQG [11] 36.47 30.25 0.15
KVQG [33] 27.18 55.38 0.13
LV2-Net [19] 29.90 92.17 0.15
[…]
Table 10: Results on the OK-VQA dataset. The results of IM-VQG are reproduced based on the official code.
Method BLEU-4 METEOR CIDEr
IM-VQG [11] 39.30 22.11 0.12
KVQG [33] 30.56 40.97 0.13
LV2-Net [19] 32.11 60.06 0.14
[…]
Table 11: Results on the A-OKVQA dataset. The results of IM-VQG are reproduced based on the official code.
We evaluated our baseline model’s generalizability in visual question generation on four datasets of natural images, namely VQG-COCO, K-VQG, OK-VQA and A-OKVQA datasets.
VQG-COCO contains 2,500 training images and 1,250 validation and test images each, designed for generating natural image-based questions. Each image is accompanied by five natural questions and five ground truth captions. Unlike other datasets, answers are not always provided.
K-VQG is a knowledge-aware VQG dataset that uniquely combines visual content with structured knowledge triplets. It is a large-scale, human-annotated dataset containing around 13,000 images, 16,000 question-answer pairs, and 6,000 knowledge triplets. Each sample consists of an image, a question, an answer, and a knowledge triplet. K-VQG allows the use of both answers or knowledge triplets as text constraints. OK-VQA represents a groundbreaking dataset in knowledge-based visual question answering, featuring 14,055 open-ended questions with 5 ground-truth answers each. Unlike previous VQA datasets that rely on pre-defined knowledge bases, OK-VQA uniquely requires models to independently retrieve and utilize external knowledge sources like Wikipedia to formulate answers. The dataset has been carefully curated through manual filtering to ensure that every question necessitates external knowledge, with deliberate reduction of frequently occurring answers to minimize dataset bias.
Building upon OK-VQA’s foundation, A-OKVQA emerges as its augmented successor, encompassing approximately 25,000 distinct questions with no overlap in question-image pairs with its predecessor. What sets A-OKVQA apart is its emphasis on comprehensive reasoning skills and diverse knowledge sources, extending beyond simple knowledge base queries. The dataset is specifically designed to challenge models with questions that demand complex commonsense reasoning about visual scenes, requiring the integration of both world knowledge and visual understanding to generate accurate answers.
Our proposed HKI-DQG consistently achieves substantial improvements over previous state-of-the-art methods across both VQG-COCO, K-VQG, OK-VQA and A-OKVQA datasets, as shown in Table 8, 9, 10 and 11. These results demonstrate the strong generalization capability of our approach and prove that using a more advanced vision-language model as a foundation can significantly enhance performance in visual question generation tasks.
E. More case studies on Diagram
Addressing the three bottlenecks in DQG task previously mentioned, Text Constraint Lost, Diagram Interpretation Bias, and Concept Ambiguity, we present additional reference cases in Figures 9, 10, and 11.
In the first case (Text Constraint Lost), as shown in Figure 9, the generated questions fail to incorporate the target text, leading to inconsistent question generation. Regarding the cause of this issue, we hypothesize that the model’s instruction processing capabilities are not fully refined, resulting in the loss of target text during the generation process.
In the second case (Diagram Interpretation Bias), as shown in Figure 10, while the generated questions assess students’ concept comprehension to some extent, they show simplified concept application compared to ground truth questions, indicating room for improvement. In response to this issue, we visualize the selected relevant background knowledge during the question generation process, as shown in Figure 10. It can be observed that these background knowledge entries do not contain any content related to the text “filter blood,” making it difficult for the model to generate questions related to this concept.
In the third case (Concept Ambiguity), as shown in Figure 11, the generated questions fail to reflect concept-related content, merely providing simple extensions of the target text. Similarly, for this issue, we visualize the selected relevant background knowledge during the question generation process, as shown in Figure 13. The questions generated in this case demonstrate basic concept awareness but fail to interact with the specific details presented in the background knowledge. This indicates that the model tends to adopt generic question patterns, lacking sufficient in-depth analysis of the specific content within the diagram, which may be related to the long-tail phenomenon present in the training samples.